zepdrix (zepdrix):

Consider the forced Van der Pol oscillator equation: \[\Large\bf\sf x''-\mu(1-x^2)x'+x\quad=\quad A \sin(\omega t)\] with \(\large \mu= 8.53\), \(\large A = 1.2\) and \(\large \omega = 2\pi/10\). \(\cancel{\text{1. Express the equation as a system of differential equations.}}\) 2. Find the fixed points of the system. 3. Linearize the system about the fixed points. 4. Determine the linear stability of the fixed points by finding the eigenvalues of the linearized system. 5. Solve the system using the 4th-order Runge-Kutta method using initial condition near and far the fixed points.

zepdrix (zepdrix):

6. Compare the numerical results with the linear stability. How far from the fixed points you need to get for the linear stability to be no longer valid. 7. Try other values of \(\mu\).

zepdrix (zepdrix):

Anyone good with MatLab? :(

zepdrix (zepdrix):

I think I've got the first part figured out, thanks to Sith.\[\Large\bf\sf x_1=x\]\[\Large\bf\sf x_2=x'\]Which gives us the system,\[\Large\bf\sf x_1'=x_2\]\[\Large\bf\sf x_2'=A \sin(\omega t) + \mu (1-x_1^2)x_2 -x_1\]

zepdrix (zepdrix):

For fixed points I set the left sides to zero I think?

zepdrix (zepdrix):

\[\Large\bf\sf 0=x_2\]\[\Large\bf\sf 0=A \sin(\omega t) + \mu (1-x_1^2)x_2 -x_1\]

zepdrix (zepdrix):

Plugging x_2 into the second equation..\[\Large\bf\sf x_1=A \sin(\omega t)\]Mmmm hopefully I'm on the right track here +_+